REASSIGNMENT OF THE 11537 cm^{−1} BAND OF HYDROGEN FLORIDE DIMER AND OBSERVATION OF THE INTERMOLECULAR COMBINATION MODE 3ν_{1} + ν_{4}

\underline{CHENG-CHI CHUANG}, SUSY N. TSANG and WILLIAM KLEMPERER, Department of Chemistry, Harvard University, Cambridge, MA 02138; HUAN-CHEN CHANG, Institute of Atomic and Molecular Sciences, Academica Sinica, P.O.Box 23-166, Taipei, Taiwan 10764, Republic of China.

We report a high resolution re-examination of K=1 subband of 3ν_{2} + ν_{6} and the 3ν_{1} + ν_{4} (K=0) combination mode of (HF)_{2} by intracavity laser induced fluorescence. The fluorescence is monitored with a germanium detector which collects only first overtone emission (2→0) of HF fragments. The band centered at 11537 cm^{−1} is shown to be K=1 of 3ν_{2} + ν_{6} [previously assigned to K=1 of ν_{1} + 2ν_{2}, J. Chem. Phys. 100, 1 (1994)]. Hydrogen interchange tunneling is almost completely quenched (Δν_{t}=0.6 GHz). Vibrational predissociation rate is rapid (4.5 GHz) but half that of 10 GHz observed in 3ν_{2}. The combination mode 3ν_{1} + ν_{4} is recorded with band origins of 11402.889(4) and 11402.868(8) cm^{−1} and rotational constants of 0.21639(17) and 0.21704(15) cm^{−1} for the two tunneling components A^{+} and B^{+}, respectively. The frequency of this intermolecular vibration, ν_{4}, upon 3ν_{1} excitation is 129.36 cm^{−1}, quite similar to that at the fundamental, suggesting only a minor dependence of the hydrogen bond vibration on the free-HF bond length. The predissociaton linewidth for 3ν_{1} + ν_{4} mode is 2.5(2) GHz, which is one order of magnitude larger than the 0.24(2) GHz of the 3ν_{1} state. The coupling of this state to the dark state of 3ν_{2} + ν_{4} + ν_{5} is suggested for the origin of the observed linewidth increase.